Experiment: Tossing two coins. Random variable đť‘‹ = number of heads a) List all possible sample space. b) Find the random variable values. c) Find the probabilities for the random variable values.



1
Expert's answer
2022-03-15T07:48:50-0400

a) Let's use the letters H and T for “head” and “tail” respectively. The sample space:

"S = \begin{Bmatrix}
 TT, HT, TH, HH\end{Bmatrix}."


b) In the case of TT (i.e we got two tails) X = 0; for HT and TH (first head, second tail or vice versa) X = 1 and for HH (both heads) X = 2.


c) The probability that we get head "p = \frac{1}{2}," that we get tail "q = 1-p=1-\frac{1}{2}=\frac{1}{2}."

The result of both tossings are independent events, for each sample point:

"P(TT)=P(T)\cdot P(T)=q\cdot q =\frac{1}{2}\cdot\frac{1}{2}=\frac{1}{4};\\
P(TH)=P(T)\cdot P(H)=q\cdot p =\frac{1}{2}\cdot\frac{1}{2}=\frac{1}{4};\\
P(HT)=P(T)\cdot P(H)=p\cdot q =\frac{1}{2}\cdot\frac{1}{2}=\frac{1}{4};\\
P(HH)=P(T)\cdot P(H)=p\cdot p =\frac{1}{2}\cdot\frac{1}{2}=\frac{1}{4}."

The probabilities for the random variable values:

"P(X=0)=P(TT)=\frac{1}{4};\\
P(X=1)=P(TH)+P(HT)=\frac{1}{4}+\frac{1}{4}=\frac{1}{2};\\
P(X=2)=P(HH)=\frac{1}{4}."